OpenStudy (dls):

The topmost and bottom most velocities of a disc are v1 and v2 in the same direction.The radius is R.Find out the velocity of the point which is R/2 above the centre?

OpenStudy (dls):

@RnR

OpenStudy (anonymous):

Pta nhi. :/ @Mashy

OpenStudy (dls):

don't be so dependent :/

OpenStudy (dls):

IIT exam mein :P @Mashy

OpenStudy (anonymous):

Rotate kr rhi hai disc?

OpenStudy (dls):

but dependency ki baat kaun kar raha h :P hmm

OpenStudy (dls):

pata nahi :/

OpenStudy (anonymous):

hmm. ho skta hai.. wait.

OpenStudy (dls):

:/

OpenStudy (dls):

 oho diagram :P

OpenStudy (anonymous):

5v1-3v2/2 ?

OpenStudy (dls):

nowhere near

OpenStudy (anonymous):

3v1-v2/4 ? calculation mistake.

OpenStudy (dls):

kaise kiya?

OpenStudy (anonymous):

v1-v+wr v2=v+wr where v is vel of centre of mass two equations. solve for v and w vel at a point r/2 v+wr/2.

OpenStudy (dls):

though the answer is 3v1"+"v2/4 ?

OpenStudy (anonymous):

ohho. haan another calc mistake. aa rha hai.

OpenStudy (dls):

v1=v+wr kahan se aati hai vise?o.O

OpenStudy (anonymous):

correction: v2=v-wr aur topmost point pe vel maximum hoti hai na.. isliye.

OpenStudy (dls):

oh :O got it :D!

OpenStudy (anonymous):

:D

OpenStudy (vincent-lyon.fr):

 If V1 and V2 are parallel, then the instant centre of rotation is on the dashed diameter of the disc, perpendicular to V1 and V2. Velocities of points along this diameter follow a simple linear law as shown.